The primary winding of an electric train transformer has 400 turns and the secondary has 50. If the input voltage is 120V(rms) what is the output voltage?


1
Expert's answer
2022-03-27T05:52:06-0400
"\frac{V_p}{V_s}=\frac{N_p}{N_s}"

"V_s=V_p\frac{N_s}{N_p}\\
=12\times \frac{50}{400}=1.5\:\rm V"